OpenStudy (anonymous):

Ari is taking a group of people to the movies. Each child costs $4 and each adult is $9. How much will he pay if the group consists of 3 adults and 4 children? A. $91 B. $52 C. $48 D. $43

OpenStudy (anonymous):

3*4+4*9 =12+36=48 babe ;)

OpenStudy (yrelhan4):

1 child costs $4 4 children will cost (cost of 1 child)*(no. of children) 1 adult costs $9 3 adults will cost (cost of 1 adult)*(no. of adults) add them to get the result..

OpenStudy (yrelhan4):

You are supposed to give a medal, btw. :)